# Quantum Yang-Mills Charges in Strongly Coupled $2D$ Lattice QCD with Three Flavors

### v1

@@ Abstract
We investigate the quantum properties of the truly gauge-invariant and conserved charges of two-dimensional Yang-Mills theories, focusing on lattice QCD in the strong coupling regime. The construction of those charges uses the integral version of the $(1+1)$-dimensional Yang-Mills equations, and they correspond to the eigenvalues of a charge operator. The gauge invariance of the charges suggests that they are not confined, hence hadronic states may carry them. Using the Feynman-Kac formalism, we evaluate the correlation functions of those charges on baryon and meson states through functional integrals over the gauge group $SU(N)$ ($N=2,3$) and Grassmannian variables — the fermionic fields. Our results show that the expectation values of the lowest non-trivial charges are nonzero for baryon and meson states but vanish for non-gauge-invariant states, supporting the interpretation that hadrons indeed carry these charges. While renormalization effects and higher-order contributions remain to be analyzed, these findings point toward a potential link between gauge-invariant charges and confinement.
@@ 1 Introduction
It is important to note that, in the strong coupling regime, the application of polymer expansions reveals significant properties. Starting with a finite lattice system in a box of volume $\Lambda$, for sufficiently small couplings $\kappa$ and $\beta$ (see [13, 14, 15]), the thermodynamic limit $\Lambda\nearrow\infty$ of correlations exists, and truncated correlations exhibit exponential tree decay. The limiting correlations are invariant under lattice translations and extend to joint analytic functions of complex coupling parameters. Additionally, our model satisfies Osterwalder-Schrader positivity [16, 17, 18, 19, 14]. This property ensures the existence of an underlying physical quantum mechanical Hilbert space $\mathcal{H}$. Within this framework, by considering translations in time and spatial coordinates, we can define self-adjoint and self-commuting energy and momentum operators on $\mathcal{H}$. Their joint spectrum is the system energy-momentum (E-M) spectrum [14, 13], which can be accessed by analyzing the analyticity properties of a lattice Fourier transform of certain correlations. A relation between Hilbert space inner products and statistical mechanical correlation of the model is provided by a Feynman-Kac type formula on the lattice [19, 15].
@@ 2.1 The classical charge operator and conserved charges
with $\kappa>0$ denoting the hopping parameter and $m>0$ the bare mass. Here $\psi$ and $\bar{\psi}$, treated as independent Grassmann variables, are nilpotents of degree 2, a consequence of the anti-commutation relations. One easily checks that the formal continuum limit $a\searrow 0$ recovers the usual smooth continuous field action.
@@ 2.3 Path integral: expectation values on the lattice
Let $\mathcal{O}(\psi,\bar{\psi},g)$, henceforth $\mathcal{O}$, be an observable. Then the inner product on $\mathcal{H}$, for $F$ and $G$ restricted to $x_{0}=1/2$, is related to the Feynmann-Kac (F-K) like formula (cf. thm. X. 68 in [28], see also [29, 19, 14]) as follows.
@@
The works of one of us and O’Carroll rigorously analyze on a lattice with finite spacing the baryonic sector of the energy-momentum spectrum in strongly coupled lattice QCD. In the 2004 study [26] (in collaboration with Schor), baryons and anti-baryons are obtained as bound states of three (anti-)quarks, with asymptotic masses $M_{s}=-3\ln\kappa+r_{s}(\kappa)$ and small mass splittings observable in $d=2$ and $d=3$, where $d$ denotes the spacetime dimensionality of the lattice. The term $r_{s}(\kappa)$ accounts for additional factors or refinements in the mass that go beyond the simple logarithmic dependence on $\kappa$, and it likely contains higher-order corrections that become relevant at specific values of $\kappa$, for small $\kappa$. While the construction of the physical Hilbert space $\mathcal{H}$ using gauge-invariant correlation functions requires a careful treatment [14], the 2004 study demonstrates that the baryonic spectrum appears as isolated energy-momentum dispersion curves within the subspace $\mathcal{H}_{o}$, generated by states with an odd number of fermions. The 2008 paper [15] extends this framework to $1+3$ dimensions, unveiling the Gell-Mann–Ne’eman eightfold way structure with baryon octets and decuplets. The masses, $M=-3\ln\kappa-3\kappa^{3}/4+\kappa^{6}r(\kappa)$, include an octet-decuplet splitting of $3\kappa^{6}/4+\mathcal{O}(\kappa^{7})$. The obtained spectral results for the one-baryon states exhibit quark confinement up to the corresponding energy threshold. As above, $r(\kappa)$ captures the contributions to the mass that arise from the $\kappa^{6}$ term in the series expansion, refining the mass prediction beyond the simple terms such as $3\ln\kappa$ and $3\kappa^{3}/4$. Here, the Hilbert space is enriched by including symmetries from both $SU(3)_{c}$ gauge and $SU(3)_{f}$ flavor groups, besides spin, and the isolated baryonic spectrum is rigorously connected to gauge-invariant states. Similar results have also been obtained for the mesonic sector of $\mathcal{H}$ [30, 31].
@@
$\expectationvalue{\mathcal{O}}_{f,g}=\frac{1}{Z_{\Lambda}}\int d\psi d\bar{\psi}d\mu(g)\,\mathcal{O}\exp\left(S_{g}+S_{f}\right),$ (2.59)
@@
where $Z_{f}$ and $\expectationvalue{\mathcal{O}}_{f}$ denote the fermionic part of the partition function and the expectation value, respectively. To be specific, $Z_{f}=\expectationvalue{1}_{f}=\int d\psi d\bar{\psi}\exp(-S_{f})$ and $\expectationvalue{\mathcal{O}}_{f}=\int d\psi d\bar{\psi}\mathcal{O}\exp(-S_{f})$. Moreover, the gauge part will be treated in a strong coupling regime that corresponds to $\exp(-S_{g})\approx 1$. Now using (2.57) and (2.62), what is to be computed for the expectation value $\expectationvalue{\mathcal{O}FG}$ should be clear. In what follows, we set [26]
@@
$\int d\mu_{C}\left(\psi,\bar{\psi}\right)\,\psi_{a\alpha f}(x)\bar{\psi}_{b\beta h}(y)=\Xi\,\delta_{ab}\delta_{\alpha\beta}\delta_{fh}\delta(x-y),\quad\mathrm{with}\quad\Xi\doteq a^{-2}\left(m+\frac{2\kappa}{a}\right)^{-1}.$ (2.64)
@@ 3.2.1 Matter part
where
@@
with
@@ 3.2.3 Total contribution: summing expectation values
In the limit $n\to\infty$—the infinite volume or thermodynamic limit [11]—the series in question along with the partition function $Z_{\Lambda}$ in the denominator diverges. Since we prefer to work on a finite nonzero lattice spacing, we do not employ the standard renormalization techniques commonly used in lattice QCD to handle such divergences. However, alternative approaches exist. In particular, the multiscale polymer expansion [13, 19] offers a systematic framework to address these issues, ensuring that divergences can be removed safely, yielding finite, well-defined results. Other regularization techniques from quantum field theory are mentioned in the remark below, which may be relevant for future investigations depending on the underlying model. For instance, Borel summation has been explored as a potential tool in this context; see [39] for related applications.
Remark 3.4.
In some lattice computations, zeta function regularization is sufficient to control divergences (for a recent application in lattice gauge theory, see [40]). Regularization systematically assigns finite values to divergent series, a standard procedure in physics (see [41]). Even in mathematics, rigorous summation methods exist, such as Borel summation [42, 43], which, for instance, assigns a finite value to certain factorial series represented as improper integrals—one well-known example being Gompertz’s constant [44]. More exotic divergent series, such as those involving factorial growth of the form $(3n)!$, can also be handled using similar techniques. This type of analysis may become particularly relevant when dealing with non-diagonal covariance matrices, cf. Isserlis-Wick theorem.
@@ 4 Conclusion
We constructed a two-dimensional Euclidean quantum gauge invariant Wilson lattice version of the gauge invariant charge using Wilson loops and developed in [5, 6, 10]. The charge is given as an infinite series. In order to understand the physical property measured by this charge and improve our knowledge about the role of local gauge invariance in quantum field models, we computed the quantum charge effects up to second order acting on non-gauge invariant states and hadrons in lattice QCD with two and three quark flavors. In the strong-coupling regime, we obtain a zero eigenvalue for the quantum charge for non-gauge invariant states and non-zero values for baryons (see 3.9) and mesons (see 3.7). In a more realistic approach, we need to incorporate higher-order contributions to the charge. In addition, renormalization effects must also be taken into account to define the lattice QCD model. Whether or not our quantum charge is a measure of confinement is a question that deserves further analysis.

### v2

@@ Abstract
We investigate the quantum properties of the truly gauge-invariant and conserved charges of two-dimensional Yang-Mills theories, focusing on lattice QCD in the strong coupling regime. The construction of those charges uses the integral version of the $(1+1)$-dimensional Yang-Mills equations, and they correspond to the eigenvalues of a charge operator. The gauge invariance of the charges suggests that they are not confined, hence hadronic states may carry them. Using the path integral formalism with imaginary time (Euclidean), we evaluate the correlation functions of those charges on baryon and meson states through functional integrals over the gauge group $SU(N)$ ($N=2,3$) and Grassmannian variables — the fermionic fields. Our results show that the expectation values of the lowest non-trivial charges are nonzero for baryon and meson states but vanish for non-gauge-invariant states, supporting the interpretation that hadrons indeed carry these charges. While renormalization effects and higher-order contributions remain to be analyzed, these findings point toward a potential link between gauge-invariant charges and confinement.
@@ 1 Introduction
Of course, there are many treatments about confinement in the lattice QCD literature, specially in the fundamental paper [11]. In a mathematically rigorous approach confinement is manifest in the exponential decay of relevant correlations, since exponential decay implies a mass gap in the energy-momentum spectrum. In the present paper we are investigating the role of the new gauge invariant conserved charges, and searching for a possible relation among those charges and confinement.
@@
We adopt the strong coupling approximation because it simplifies the type of contributions to be considered when computing correlations (Green’s functions) and, most important, it makes the evaluation of the gauge group integrals with the Haar measure much simpler. In addition, we are interested in evaluating the expectation values of the charges on bound states, such as hadronic particle fields. It is worthwhile mentioning that the strong coupling regime has shown to be a good approximation. In previous papers, by one of us, the existence of hadronic particles and some of their bound states was rigorously confirmed and quark confinement was verified up to some energy threshold in the energy-momentum spectrum [14, 15]. We also chose to work in $(1+1)$-dimensions since the spin is not really present, simplifying the algebraic computations. The strong coupling and the dimensionality are not mandatory.
It is important to note that, in the strong coupling regime, the application of polymer expansions reveals significant properties. Starting with a finite lattice system in a box of volume $\Lambda$, for sufficiently small couplings $\kappa$ and $\beta$ (see [16, 17, 18]), the thermodynamic limit $\Lambda\nearrow\infty$ of correlations exists, and truncated correlations exhibit exponential tree decay. The limiting correlations are invariant under lattice translations and extend to joint analytic functions of complex coupling parameters. Additionally, our model satisfies Osterwalder-Schrader positivity [19, 20, 21, 22, 17]. This property ensures the existence of an underlying physical quantum mechanical Hilbert space $\mathcal{H}$. Within this framework, by considering translations in time and spatial coordinates, we can define self-adjoint and self-commuting energy and momentum operators on $\mathcal{H}$. Their joint spectrum is the system energy-momentum (E-M) spectrum [17, 16], which can be accessed by analyzing the analyticity properties of a lattice Fourier transform of certain correlations. A relation between Hilbert space inner products and statistical mechanical correlation of the model is provided by a path integral (Feynman-Kac) type formula on the lattice [22, 18].
@@ 2.1 The classical charge operator and conserved charges
with $\kappa>0$ denoting the hopping parameter and $m>0$ the bare mass. Here, $\psi_{b\beta f}$ and ${\bar{\psi}}_{b\beta f}$ are Wilson fermions (no doubling), with $b=1,2$, labelling the color $SU(2)$ doublet representation components, $\beta$ is the spin state index (particle and anti-particle), and $f$ being the flavor (isospin) label. In addition, $\psi$ and $\bar{\psi}$, treated as independent Grassmann variables, are nilpotents of degree 2, a consequence of the anti-commutation relations. One easily checks that the formal continuum limit $a\searrow 0$ recovers the usual smooth continuous field action.
@@ 2.3 Path integral: expectation values on the lattice
Let $\mathcal{O}(\psi,\bar{\psi},g)$, henceforth $\mathcal{O}$, be an observable. Then the inner product on $\mathcal{H}$, for $F$ and $G$ restricted to $x_{0}=1/2$, is related to the Feynman-Kac (F-K) like formula (cf. thm. X. 68 in [31], see also [32, 22, 17]) as follows.
@@
The particle spectrum was analyzed in Refs. [33, 34] at an early stage of lattice QCD, in the strong coupling approximation. Some works by one of us and O’Carroll extends these analyses to a rigorous level Using a lattice with fixed finite spacing, the baryonic sector of the energy-momentum spectrum in strongly coupled lattice QCD. In the 2004 study [29] (in collaboration with Schor), baryons and anti-baryons are obtained as bound states of three (anti-)quarks, with asymptotic masses $M_{s}=-3\ln\kappa+r_{s}(\kappa)$ and small mass splittings observable in $d=2$ and $d=3$, where $d$ denotes the spacetime dimensionality of the lattice. The term $r_{s}(\kappa)$ accounts for additional factors or refinements in the mass that go beyond the simple logarithmic dependence on $\kappa$, and it likely contains higher-order corrections that become relevant at specific values of $\kappa$, for small $\kappa$. While the construction of the physical Hilbert space $\mathcal{H}$ using gauge-invariant correlation functions requires a careful treatment [17], the 2004 study demonstrates that the baryonic spectrum appears as isolated energy-momentum dispersion curves within the subspace $\mathcal{H}_{o}$, generated by states with an odd number of fermions. The 2008 paper [18] extends this framework to $1+3$ dimensions, unveiling the Gell-Mann–Ne’eman eightfold way structure with baryon octets and decuplets. The masses, $M=-3\ln\kappa-3\kappa^{3}/4+\kappa^{6}r(\kappa)$, include an octet-decuplet splitting of $3\kappa^{6}/4+\mathcal{O}(\kappa^{7})$. The obtained spectral results for the one-baryon states exhibit quark confinement up to the corresponding energy threshold. As above, $r(\kappa)$ captures the contributions to the mass that arise from the $\kappa^{6}$ term in the series expansion, refining the mass prediction beyond the simple terms such as $3\ln\kappa$ and $3\kappa^{3}/4$. Here, the Hilbert space is enriched by including symmetries from both $SU(3)_{c}$ gauge and $SU(3)_{f}$ flavor groups, besides spin, and the isolated baryonic spectrum is rigorously connected to gauge-invariant states. Similar results have also been obtained for the mesonic sector of $\mathcal{H}$ [35, 36]. Two-particle baryonic bound states such as the deuterum were obtained e.g. in Refs. [14, 15].
@@
$\expectationvalue{\mathcal{O}}_{f,g}=\frac{1}{Z_{\Lambda}}\int d\psi d\bar{\psi}d\mu(g)\,\mathcal{O}\exp\left(-S_{g}-S_{f}\right),$ (2.59)
@@
where $Z_{f}$ and $\expectationvalue{\mathcal{O}}_{f}$ denote the fermionic part of the partition function and the expectation value, respectively. To be specific,
$Z_{f}=\int d\psi d\bar{\psi}\exp(-S_{f})$ (2.63)
and
$\expectationvalue{\mathcal{O}}_{f}=\frac{1}{Z_{f}}\,\int d\psi d\bar{\psi}\mathcal{O}\exp(-S_{f})$ (2.64)
Moreover, the gauge part will be treated in a strong coupling regime that corresponds to $\exp(-S_{g})\approx 1$. Now using (2.57) and (2.62), what is to be computed for the expectation value $\expectationvalue{\mathcal{O}FG}$ should be clear. In what follows, we set [29]
@@
$\int d\mu_{C}\left(\psi,\bar{\psi}\right)\,\psi_{a\alpha f}(x)\bar{\psi}_{b\beta h}(y)=\Xi\,\delta_{ab}\delta_{\alpha\beta}\delta_{fh}\delta(x-y),$ (2.66)
with
$\Xi\doteq a^{-2}\left(m+\frac{2\kappa}{a}\right)^{-1}.$ (2.67)
@@ 3.2.1 Matter part
where $Z_{f}$ and $\Xi$ are defined respectively in (2.63) and (2.67), and $E_{f}^{\pm}$, for mesons, is given by
@@
recalling that $Z_{f}$ and $\Xi$ are defined respectively in (2.63) and (2.67). Here $E_{f}^{\pm}$, for baryons, is given by
@@ 3.2.3 Total contribution: summing expectation values
In the limit $n\to\infty$—the infinite volume or thermodynamic limit [12]—the series in question along with the partition function $Z_{\Lambda}$ in the denominator diverges. Since we prefer to work on a finite nonzero lattice spacing, we do not employ the standard renormalization techniques commonly used in lattice QCD to handle such divergences. However, alternative approaches exist. In particular, the multiscale polymer expansion [16, 22] offers a systematic framework to address these issues, ensuring that divergences can be removed safely, yielding finite, well-defined results. Other regularization techniques from quantum field theory can be used, and may be relevant for future investigations depending on the underlying model.
The average appearing in (3.15) are computed and given in Appendix B for hadrons, in the case of lattice QCD models with two and three flavors. Note that our explicit computations are in agreement with the flavor symmetry of the model. By considering the hadronic particles wave functions in terms of quarks and anti-quarks, and their linear combinations, the values of the quantum charges of the hadrons follow. For instance, in the three flavor case and considering the pre-factors given in (3.7), the charge values for $\pi^{+}$ and $\pi^{-}$ coincide, and it vanishes for $\pi^{0}$. Regarding baryons and using (3.9), we obtain that the charge value for $\Delta^{++}$ differs from $p$ and $n$ (proton and neutron), which are equal.
The reason for the vanishing charge of $\pi^{0}$ is algebraic, and is due to the fact that its wave function is the linear combination $\pi^{0}\equiv\left(u\,{\bar{u}}-d\,{\bar{d}}\right)/\sqrt{2}$, whilst for $\pi^{\pm}$ the linear combination with a minus sign is not there, as $\pi^{+}\equiv u\,{\bar{d}}$, and $\pi^{-}\equiv d\,{\bar{u}}$. Indeed, using (3.7) and the tables in Appendix B gives non-zero charge for $\pi^{\pm}$ and a vanishing charge for $\pi^{0}$.
We also considered gauge variant states for which we obtain vanishing charges.
@@ 4 Conclusion
We constructed a two-dimensional Euclidean quantum gauge invariant Wilson lattice version of the gauge invariant charge using Wilson loops and developed in [5, 6, 10]. The charge is given as an infinite series. In order to understand the physical property measured by this charge and improve our knowledge about the role of local gauge invariance in quantum field models, we computed the quantum charge effects up to second order acting on non-gauge invariant states and hadrons in lattice QCD with two and three quark flavors. In the strong-coupling regime, we obtain a zero eigenvalue for the quantum charge for non-gauge invariant states and non-zero values for baryons (see (3.9)) and mesons (see (3.7)). In a more realistic approach, we need to incorporate higher-order contributions to the charge. In addition, renormalization effects must also be taken into account to define the lattice QCD model. Whether or not our quantum charge is a measure of confinement is a question that deserves further analysis.
Of course it would be very desirable to understand the role of the charge $Q_{M}^{(2)}$ in a more phenomenological way, e.g. analyzing some standard decay processes as $\pi^{0}\rightarrow 2\,\gamma$, $\pi^{+}\rightarrow\mu^{+}+\nu_{\mu}$, and $\Delta^{++}\rightarrow p+\pi^{+}$. There is a wide open pathway to explore on those lines.
In addition, it would also be interesting to extend our analysis to the quantum version of the higher order charges $Q_{M}^{(j)}$, given in (2.41). And, more importantly, to extend the analysis to the case of real QCD in $(3+1)$-dimensions, and verify if the real hadrons carry the quantum version of the charges constructed in [9].